OpenStudy (anonymous):

PLEASE HELP I NEED TO KNOW how do you do the 30-60-90 in geometry with triangles

OpenStudy (kainui):

Well, depends on what you want to do, but you know the Pythagorean theorem right? a^2+b^2=c^2 and c is the hypotenuse while a and b are the legs that meet at the right angle:  So you can reflect it over to get a perfect equilateral triangle, this lets you do some tricks like say that side c=2*a so you can plug that into the pythagorean theorem: \[\Large a^2+b^2=(2*a)^2 \\ \Large a^2+b^2 = 4*a^2\] subtract one of the a^2 from both sides so the left side only has a b^2 left and we went from 4 a^2 down to 3 of the a^2 terms so we have: \[\Large b^2 = 3a^2\] Take the square root of both sides: \[\Large b= a \sqrt{3}\] So there you have a pretty simple algebra rule for playing with 30-60-90 triangles.

OpenStudy (anonymous):

thanks that helps a little

OpenStudy (kainui):

Yeah if you have specific questions I can help you but I mean this is pretty general haha.

OpenStudy (anonymous):

she gave us a tutorial example but it doesn't help that much